# Calibration in Deep Learning: A Survey of the State-of-the-Art

### v3

@@ 1 Introduction
Figure 1: An illustration of model calibration. Uncalibrated model (left) that trained with standard cross-entropy loss and calibrated model (right) that trained with focal loss ($\gamma=5$), have similar predictive performance on a binary classification task (accuracy is 83.8% and 83.4% respectively), but the right one is better calibrated. Top: The reliability diagram plots with 10 bins. The diagonal dash line presents perfect calibration (on a specific bin, confidence is equal to accuracy.), Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) are used to measure model calibration performance. Bottom: the posterior distribution of the two models on 1000 samples, the one from calibrated model is better distributed.
@@ 2.1 Definitions
In classification tasks, for a given input variable $X$ and a categorical variable $Y\in\{1,2,...,k\}$, assume we have a neural network model $f$ which maps input variable $\mathbf{x}$ to a categorical distribution $p=\{p_{1},...,p_{k}\}$ over $k$ classes $\{y_{1},...,y_{k}\}$: $f:D\rightarrow\Delta$, where $\Delta$ is the $k-1$ dimensional standard probability simplex and $\Delta=\{p\in[0,1]^{k}|\sum_{i=1}^{k}p_{i}=1\}$. Calibration measures the degree of the match between predicted probability $p$ and the true correctness likelihood. A model $f$ is perfectly calibrated on if and only if:
@@
Instead of probability distribution, the argmax calibration (?, ?, ?) takes only the maximum probability into consideration:
$\mathbb{P}(Y\in\arg\max(p)|\max(f(X))=p^{*})=p^{*}$ (2)
@@ 2.2.1 Model Complexity
While increasing the depth and width of neural networks helps to obtain highly predictive power, it also negatively increases calibration errors. Empirical evidence has shown that this poor calibration is linked to overfitting on the negative log-likelihood (NLL) (?, ?). The over-parameteriaztion is one of the main causes of overfitting. Concretely, along with the standard NLL-based model training, when classification error is minimized, keeping training will further push the model to minimize NLL on the training data, i.e., push the predicted softmax probability distribution as close as possible to the ground-truth distribution (which is usually one-hot representation). Model overfitting starts by exhibiting increased test NLL, and then the model becomes overconfident (?). This phenomena becomes more severe when training a high capacity model with limited model.
@@ 2.2.2 Data Issues
Another important aspects that directly impact model calibration are: data quantity, i.e., the scale of dataset and data quality such as relevance, consistency, completeness, etc. Training high-capacity (over-parameterized) networks with scarce data can easily cause overfitting and have an overconfident model. Data augmentation is an effective way to alleviate this phenomenon and brings implicit calibration effects (?, ?). The recent pretrain-finetun paradigm offers the possibility of reducing overfitting caused by limited data (?). Besides having a decent amount of training data, we should also highlight the importance of addressing the following issues:
  • Data noise: this refers to the presence of erroneous or misleading information in datasets, which can adversely affect model training and performance. Noisy data can arise from various sources such as systematic or measurement errors when pulling historical data from production models, labeling or annotation mistakes caused by human annotators, or outliers  (?). Those factors can lead to biased model estimates, reduced predictive accuracy, and poor generalization to unseen data. This contributes to model miscalibration. Some methods has been proposed to mitigate the impact of data noise on model calibration such as robust loss functions or adversarial training, have shown promise in improving model robustness to noisy data (?, ?). Also Ensemble methods have been proofed effective in identifying and mitigating the effects of noisy data on model performance (?, ?).
  • Data imbalance. It can significantly impact model calibration. When a dataset is imbalanced, meaning that one class or outcome is significantly more prevalent than others, it can lead to biased predictions and poor calibration. This is because the model may become overly confident in predicting the majority class while being less certain about the minority classes. Focal loss  (?, ?) has shown its effectiveness in calibrating models that trained with imbalanced datasets (?).
  • Data biases and fairness. Extensive studies have highlighted the presence of biases in datasets. Biases arises due to historical disparities, societal prejudices, or systemic inequalities. For instance, facial recognition datasets often exhibit racial and gender biases (?). Mitigating bias and ensuring fairness are imperative steps towards building models that exhibit better calibration and reliability (?, ?). This involves a multifaceted approach that begins with critically examining the data collection process to identify and rectify sources of bias. Techniques such as data augmentation (?, ?), bias mitigation algorithms (?), and fairness-aware model training (?) can help mitigate biases during the model development stage.
@@ 3 Calibration Measurements
3.0.1 Expected Calibration Error (ECE)
@@
3.0.2 Maximum Calibration Error (MCE)
@@
3.0.3 Classwise ECE (CECE)
@@ 3.0.4 Adaptive ECE (AECE)
For a perfectly calibrated classifier, those calibration erros should equal 00.
3.0.5 Reliability Diagram
@@ 4 Calibration Methods
for tree=draw, semithick, rounded corners, edge = draw, semithick, anchor = east, grow = east, forked edge, s sep = 2mm, l sep = 2mm, fork sep = 2mm, [Calibration [Large Language Models [Contextual Calibration (?)] [Prototypical Calibration (?)] [Domain-Context Calibration (?)] [Batch Calibration (?)] ] [Hybird Methods [LS + Mixup (?)] [Ensemble TS (ETS) (?)] [TS + MC dropout (?)] ] [ Uncertainty Estimation [Bayesian Methods [Finite-State Probabilistic RNNs (?)] [MC Dropout (?)] [Bayesian RNNs (?)] [BNNs (?)] ] [Ensemble Methods [Deep Ensembles (?)] ] ] [Regularization Methods [Explicit Regularization [Huber loss (?)] [DCA (?)] [Entropy Regularization (?)] ] [Implicit Regularization [Focal Loss (?)] [FLSD (?)] [Label Smooth(LS) (?)] [Mixup (?)] ] [Differentiable Proxy [MMCE (?)] [Meta-Calibration (?)] [AvUC (?)] [MDCA (?)] ] ] [Post-hoc Methods [Local TS (LTS) (?)] [Dirichlet calibration (?)] [Bin-wise TS (BTS) (?)] [Attended TS (ATS) (?)] [Temperature Scaling (TS) (?)] ] ]
Figure 3: The categorization of calibration methods and the representative methods in each category. The regularization methods are further divided into explicit, implicit regularization and differentiable calibration proxies.
@@ 4.1.1 Temperature Scaling
where $\tau>0$ for all classes is used as a scaling factor to soften model predicted probability, it controls the model’s confidence by adjusting the sharpness of distribution so that the model prediction is not too certain (overconfident) or too uncertain (underconfident). The optimal temperature value is obtained by minimizing negative log likelihood loss (NLL) on the validation dataset.:
@@ 4.1.2 Temperature Scaling Extensions
The goal of post-hoc calibration on a validation dataset is to learn a calibration map (also known as the canonical calibration function of probability (?)) which transforms uncalibrated probabilities into calibrated ones. Many TS extensions aim to find a proper calibration map.   Kull et al. (?) proposed Dirichlet calibration which assumes probability distributions are parameterized Dirichlet distributions:
@@
where $\alpha^{j}=\{\alpha^{j}_{1},\alpha^{j}_{2},...,\alpha^{j}_{k}\}$ are Dirichlet parameters for $j$-th class. The proposed Dirichlet calibration map family coincides with the Beta calibration family (?). Besides, it provides uniqueness and interpretability as compared to generic canonical parametrization.    (?) suggested that TS has difficulties in finding optimal $\tau$ when the validation set has a limited number of samples. They proposed attended temperature scaling (ATS) to alleviate this issue by increasing the number of samples in validation set. The key idea is to gather samples from each class distribution. Let’s assume $p(y|x)$ is the predicted probability. ATS first divides the validation set into $K$ subsets with $p(y=k|x),k\in[1,K]$, which allows to add more $y\neq k$ samples using the Bayeisan Theorem (?) as the selection criterion:
@@
Remarks: Although using a single global hyper-parameter, TS remains a popular approach due to its effectiveness and accuracy-preserving (?). Post-hoc calibration usually works well without a huge amount of validation data and is thus data efficient. The calibration procedure is decoupled from training and does not introduce training complexity. On the other hand, post-hoc calibration is less expressive and suffers difficulty in approximating the canonical calibration function when data is not enough (?).
@@ 4.2.1 Explicit Regularization
$\mathcal{L}(\theta)=-\sum\log p(y|x)-\alpha H(p(y|x)).$ (15)
directly penalizes predictive distributions that have low entropy, prevents these peaked distributions, and ensures better model generalization. Recently, (?) analyzed some current calibration methods such as label smoothing (?), focal loss (?), and explicit confidence penalty (?) from a constrained-optimization perspective and pointed out those can be seen as approximations of a liner penalty. The authors proposed to add a margin $m_{a}$ into learning objective:
$\min\mathcal{L}_{CE}+\alpha\sum_{k}\max(0,\max_{j}(g_{j})-j_{k}-m_{a})$ (16)
(?) proposed to directly add the difference between confidence and accuracy (DCA) to standard cross-entropy as a penalty term to improve calibration. (?) further extend this non-differentiable term to a differentiable version using Huber loss, which is used as an explicit regularisation. They also proposed to use a hyper-parameter to control the transition of explicit L1 and L2 regularizers. The Huber loss is then combined with focal loss which provides implicit regularisation effects.
@@ 4.2.2 Implicit Regularization: Focal Loss and Its Extensions
Focal loss (?) was originally proposed to alleviate the class imbalance issue in object detection:$\mathcal{L}_{f}=-\sum_{i=1}^{N}(1-p_{i})^{\gamma}\log p_{i}$ where $\gamma$ is a hyperparameter. It has been recently shown that focal loss can be interpreted as a trade-off between minimizing Kullback–Leibler (KL) divergence and maximizing the entropy, depending on $\gamma$ (?):
@@
The first term pushes the model to learn a probability $p$ to have a high value (confident, as close as possible to ground-truth label distribution, which is usually one-hot representation). The second term is constant. The last term regularizes probability to not be too high (overconfident).   Mukhoti et al. (?) empirically observed that $\gamma$ plays an important role in implicitly regularizing entropy and weights. However, finding an appropriate $\gamma$ is challenging for all samples in the datasets. Thus, they proposed sample-dependent scheduled $\gamma$ (FLSD) based on the Lambert-W function (?). They have shown that scheduling $\gamma$ values according to different confidence ranges helps to improve model calibration on both in-domain and out-of-domain (OOD) data.

### v4

@@ 1 Introduction
(a) Reliability diagram (uncalibrated)
(b) Predictive probability (uncalibrated)
(c) Reliability diagram (calibrated)
(d) Predictive probability (calibrated)
Figure 1: Comparison of reliability diagrams and predictive probability between uncalibrated and calibrated binary classification models. (a) and (b) correspond to the uncalibrated model trained with standard cross-entropy loss; (c) and (d) show the calibrated model trained with focal loss ($\gamma=5$). Both models achieve similar accuracy (83.8% and 83.4%, respectively), but differ in calibration. Reliability diagrams (a) and (c) plot predicted confidence (x-axis) versus empirical accuracy (y-axis) using 10 bins. The dashed diagonal represents perfect calibration, while the Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) quantify calibration performance. Histograms (b) and (d) display the distribution of predicted probabilities on 1000 test samples. Pink bars indicate the calibration error (i.e., the absolute difference between average confidence and empirical accuracy) in each bin.
@@ 2.1 Definitions
In classification tasks, for a given input variable $X$ and a categorical variable $Y\in\{1,2,...,k\}$, assume we have a neural network model $f$ which maps input variable $\mathbf{x}$ to a categorical distribution $p=\{p_{1},...,p_{k}\}$ over $k$ classes $\{y_{1},...,y_{k}\}$: $f:D\rightarrow\Delta$, where $\Delta$ is the k1k-1 dimensional standard probability simplex and $\Delta=\{p\in[0,1]^{k}|\sum_{i=1}^{k}p_{i}=1\}$. Calibration measures the degree of the match between predicted probability $p$ and the true correctness likelihood. A model $f$ is perfectly calibrated on if and only if:
@@
Instead of using the full probability distribution, the argmax calibration (?, ?, ?) takes only the maximum probability into consideration:
$\mathbb{P}(Y=\arg\max(p)|\max(f(X))=p^{*})=p^{*}$ (2)
@@ 2.2.1 Model Complexity
While increasing the depth and width of neural networks helps to obtain highly predictive power, it also negatively increases calibration errors. Empirical evidence has shown that this poor calibration is linked to overfitting on the negative log-likelihood (NLL) (?, ?). The over-parameterization is one of the main causes of overfitting. Concretely, along with the standard NLL-based model training, when classification error is minimized, keeping training will further push the model to minimize NLL on the training data, i.e., push the predicted softmax probability distribution as close as possible to the ground-truth distribution (which is usually one-hot representation). Model overfitting starts by exhibiting increased test NLL, and then the model becomes overconfident (?). This phenomena becomes more severe when training a high capacity model with limited data.
@@ 2.2.2 Data Issues
Another important aspects that directly impact model calibration are: data quantity, i.e., the scale of dataset and data quality such as relevance, consistency, completeness, etc. Training high-capacity (over-parameterized) networks with scarce data can easily cause overfitting and have an overconfident model. Data augmentation is an effective way to alleviate this phenomenon and brings implicit calibration effects (?, ?). The recent pretrain-finetune paradigm offers the possibility of reducing overfitting caused by limited data (?). Besides having a decent amount of training data, we should also highlight the importance of addressing the following issues:
  • Data noise: This refers to the presence of erroneous or misleading information in datasets, which can adversely affect model training and performance. Noisy data can arise from various sources such as systematic or measurement errors when pulling historical data from production models, labeling or annotation mistakes caused by human annotators, or outliers  (?). Those factors can lead to biased model estimates, reduced predictive accuracy, and poor generalization to unseen data. This contributes to model miscalibration. Some methods have been proposed to mitigate the impact of data noise on model calibration such as robust loss functions or adversarial training, have shown promise in improving model robustness to noisy data (?, ?). Ensemble methods have also been proven effective in identifying and mitigating the effects of noisy data on model performance (?, ?)..
  • Data imbalance: It can significantly impact model calibration. When a dataset is imbalanced, meaning that one class or outcome is significantly more prevalent than others, it can lead to biased predictions and poor calibration. This is because the model may become overly confident in predicting the majority class while being less certain about the minority classes. Focal loss  (?, ?) has shown its effectiveness in calibrating models that trained with imbalanced datasets (?).
  • Data biases and fairness: Extensive studies have highlighted the presence of biases in datasets. Biases arises due to historical disparities, societal prejudices, or systemic inequalities. For instance, facial recognition datasets often exhibit racial and gender biases (?). Mitigating bias and ensuring fairness are imperative steps towards building models that exhibit better calibration and reliability (?, ?). This involves a multifaceted approach that begins with critically examining the data collection process to identify and rectify sources of bias. Techniques such as data augmentation (?, ?), bias mitigation algorithms (?), and fairness-aware model training (?) can help mitigate biases during the model development stage.
@@ 3 Calibration Measurements
3.1 Expected Calibration Error (ECE)
@@
3.2 Maximum Calibration Error (MCE)
@@
3.3 Brier Score
The Brier Score (?) is a proper scoring rule that measures both discrimination and calibration of probabilistic predictions. For binary classification, it computes the mean squared error between predicted probabilities and actual outcomes. The metric can be formulated as
$\textsc{BS}=\frac{1}{N}\sum_{i=1}^{N}(f_{i}-y_{i})^{2}$ (6)
where $f_{i}$ represents the predicted probability for the positive class and $y_{i}\in{0,1}$ is the true binary label. For multi-class problems with $K$ classes, the Brier Score can be extended as
$\textsc{BS}=\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{K}(f_{i,c}-y_{i,c})^{2}$ (7)
where $f_{i,c}$ is the predicted probability for class $c$ and $y_{i,c}$ is the binary indicator for the true class.
The Brier Score ranges from 0 to 1, where 0 indicates perfect calibration and discrimination. It can be decomposed into reliability (calibration), resolution, and uncertainty components, providing insights into different aspects of predictive performance.
3.4 Classwise ECE (CECE)
@@ 3.5 Adaptive ECE (AECE)
For a perfectly calibrated classifier, those calibration errors should equal 0.
3.6 Reliability Diagram
@@ 4 Calibration Methods
for tree= draw, semithick, rounded corners, drop shadow, text centered, edge=draw, semithick, anchor=east, grow=east, forked edge, s sep=2mm, l sep=2mm, fork sep=2mm [Calibration, top color=red!20, bottom color=red!40 [Large Language Models, top color=blue!20, bottom color=blue!40, for descendants=fill=blue!10 [Contextual Calibration (?)] [Prototypical Calibration (?)] [Domain-Context Calibration (?)] [Batch Calibration (?)] [PriDe (?)] [Thermometer (?)] [Atomic Calibration(?)] ] [Hybrid Methods, top color=green!20, bottom color=green!40, for descendants=fill=green!10 [LS + Mixup (?)] [Ensemble TS (ETS) (?)] [TS + MC Dropout (?)] [Conformal Prediction + MC Dropout (?)] ] [Uncertainty Estimation, top color=purple!20, bottom color=purple!40, for descendants=fill=purple!10 [Sample-weighted Gradient [BSCE‑GRA (?)] ] [Bayesian Methods [Finite-State Probabilistic RNNs (?)] [MC Dropout (?)] [Bayesian RNNs (?)] [BNNs (?)] ] [Ensemble Methods [Deep Ensembles (?)] [Adaptive Calibrator Ensemble (?)] ] ] [Regularization Methods, top color=orange!20, bottom color=orange!40, for descendants=fill=orange!10 [Explicit Regularization [Huber Loss (?)] [DCA (?)] [Entropy Regularization (?)] ] [Implicit Regularization [Focal Loss (?)] [FLSD (?)] [Label Smoothing (LS) (?)] [Mixup (?)] ] [Differentiable Proxy [MMCE (?)] [Meta-Calibration (?)] [AvUC (?)] [MDCA (?)] ] ] [Post-hoc Methods, top color=gray!20, bottom color=gray!40, for descendants=fill=gray!10 [Top-versus-All (TvA) (?)] [Density‑Aware Calibration (DAC) (?)] [One-vs-All Surrogate Loss (?)] [Local TS (LTS) (?)] [Dirichlet Calibration (?)] [Bin-wise TS (BTS) (?)] [Attended TS (ATS) (?)] [Temperature Scaling (TS) (?)] ] ]
Figure 3: Categorization of calibration methods and representative approaches.
@@ 4.1.1 Temperature Scaling
where $g_{i}$ is the logit and $\tau>0$ for all classes is used as a scaling factor to soften model predicted probability, it controls the model’s confidence by adjusting the sharpness of distribution so that the model prediction is not too certain (overconfident) or too uncertain (underconfident). The optimal temperature value is obtained by minimizing negative log likelihood loss (NLL) on the validation dataset.:
@@ 4.1.2 Temperature Scaling Extensions
The goal of post-hoc calibration on a validation dataset is to learn a calibration map (also known as the canonical calibration function of probability (?)) which transforms uncalibrated probabilities into calibrated ones. Many TS extensions aim to find a proper calibration map.
Kull et al. (?) proposed Dirichlet calibration which assumes probability distributions are parameterized Dirichlet distributions:
@@
where $\alpha^{j}=\{\alpha^{j}_{1},\alpha^{j}_{2},...,\alpha^{j}_{k}\}$ are Dirichlet parameters for the $j$-th class among $k$ total classes. The proposed Dirichlet calibration map family coincides with the Beta calibration family (?). Besides, it provides uniqueness and interpretability as compared to generic canonical parametrization.
Mozafari et al. (?) suggested that TS has difficulties in finding optimal $\tau$ when the validation set has a limited number of samples. They proposed attended temperature scaling (ATS) to alleviate this issue by increasing the number of samples in validation set. The key idea is to gather samples from each class distribution. Let’s assume $p(y|x)$ is the predicted probability. ATS first divides the validation set into $K$ subsets with $p(y=k|x),k\in[1,K]$, which allows to add more $y\neq k$ samples using the Bayesian Theorem (?) as the selection criterion:
@@
Verma and Nalisnick (?) revisit the “learning to defer” (L2D) framework, which allows a model to defer decisions to a human expert when appropriate. They identify that the prevalent multiclass L2D approach by  (?), which uses a single softmax over classes plus a reject option, is intrinsically miscalibrated with respect to predicting expert correctness—its softmax parameterization can yield probabilities summing to more than one and does not reflect true likelihoods of expert success. To correct this, the authors propose an alternate scheme using one‑vs‑all (OvA) classifiers: each class (including expert correctness) is modeled with its own sigmoid output, forming a vector of calibrated probabilities. They derive a consistent surrogate loss (denoted $\psi_{\text{OvA}}$) that reduces to multiple binary classification problems. This construction guarantees well-calibrated probabilities while maintaining consistency with the 0–1 L2D objective. Empirical evaluation across tasks such as hate speech detection, galaxy classification, and skin lesion diagnosis shows that the OvA‑based model attains calibration without sacrificing—often improving—accuracy compared to the softmax baseline.
Le Coz et al. (?) proposed the Top-versus-All (TvA) post‑hoc calibration method, which reformulates multiclass confidence calibration into a single binary task. They construct a surrogate classifier by taking the maximum softmax score and assigning a binary label of correctness. This setup allows applying standard binary calibration tools such as Temperature Scaling (?), Histogram Binning (?) and Dirichlet Calibration (?) more effectively, avoiding issues of multiclass imbalance and complexity. In the case of scaling methods, they replace the usual cross‑entropy with binary cross‑entropy, creating stronger gradients that penalize incorrect overconfidence. They also introduce simple $L_{2}$ regularization for Vector/Dirichlet Scaling to reduce overfitting in large‑class regimes. Extensive experiments on image and text–based models (including PLMs and LLMs) demonstrate that TvA improves calibration metrics (like ECE), preserves predicted labels, and is easy to plug into existing pipelines.
Tomaniet al. (?) introduced Density‑Aware Calibration (DAC), a post‑hoc method designed to improve calibration under domain shifts and OOD scenarios. DAC uses k‑nearest neighbor density estimates in hidden-layer feature space to measure how “in‑distribution” a test sample is, then rescales its logits based on this density before applying any standard post‑hoc calibrator $h$ (e.g., TS, Dirichlet/Vector Scaling, splines). This preserves in-domain accuracy while significantly reducing Expected Calibration Error (ECE) under corruptions and OOD. Experiments on CIFAR‑C and ImageNet‑C with CNNs and Transformers show that DAC enhances robustness and calibration without retraining the original model.
4.1.3 Practical Considerations
Despite using only a single global hyperparameter, TS remains widely adopted due to its simplicity, effectiveness, and ability to preserve accuracy (?). Post-hoc methods are appealing because they decouple calibration from training, require minimal computational overhead, and perform well even with limited validation data. However, their expressiveness is often constrained. While extensions like Dirichlet (?), Bin-wise (?), and Local TS (?) offer more flexibility, they introduce complexity and may require domain-specific tuning. Some methods address data sparsity directly, such as Attended TS (?), which enriches calibration samples through class-aware sampling.
Recent advances extend post-hoc calibration to new settings. Verma and Nalisnick (?) improve deferral calibration using one-vs-all classifiers, while Le Coz et al. (?) reduce multiclass calibration to a binary task via Top-versus-All scoring. Tomani et al. (?) introduce density-aware scaling to improve calibration under distribution shift. Nonetheless, post-hoc methods depend heavily on the quality and representativeness of validation data. Their limited capacity to approximate complex calibration maps may reduce effectiveness under domain shift or when applied to more nuanced prediction tasks (?).
@@ 4.2.1 Explicit Regularization
$\mathcal{L}(\theta)=-\sum\log p(y\mid x)-\alpha H(p(y\mid x))$ (17)
directly penalizes predictive distributions with low entropy, discouraging overly confident (peaked) outputs and promoting better generalization.
Recently, (?) analyzed popular calibration methods such as label smoothing (?), focal loss (?), and confidence penalties (?) from a constrained optimization perspective. They showed that these methods can be interpreted as approximations of a linear penalty. Based on this analysis, the authors proposed adding a margin $m_{a}$ to the learning objective:
$\min\mathcal{L}_{CE}+\alpha\sum_{k}\max\bigl{(}0,\max_{j}(g_{j}^{(k)})-g_{y_{k}}^{(k)}-m_{a}\bigr{)}$ (18)
where $g_{j}^{(k)}$ denotes the logit for class $j$ for input $x_{k}$, and $g_{y_{k}}^{(k)}$ is the logit corresponding to the ground-truth class $y_{k}$.
Liang et al., (?) proposed to directly add the difference between confidence and accuracy (DCA) to standard cross-entropy as a penalty term to improve calibration. (?) further extend this non-differentiable term to a differentiable version using Huber loss, which is used as an explicit regularization. They also proposed to use a hyper-parameter to control the transition of explicit L1 and L2 regularizer. The Huber loss is then combined with focal loss which provides implicit regularization effects.
@@ 4.2.2 Implicit Regularization: Focal Loss and Its Extensions
[…]
@@
The first term pushes the model to learn a probability $p$ to have a high value (confident, as close as possible to ground-truth label distribution, which is usually one-hot representation). The second term is constant. The last term regularizes probability to not be too high (overconfident).
Mukhoti et al. (?) empirically observed that $\gamma$ plays an important role in implicitly regularizing entropy and weights. However, finding an appropriate $\gamma$ is challenging for all samples in the datasets. Thus, they proposed sample-dependent scheduled $\gamma$ (FLSD) based on the Lambert-W function (?). They have shown that scheduling $\gamma$ values according to different confidence ranges helps to improve model calibration on both in-domain and out-of-domain (OOD) data.